value of cos(-660)*tan(510)
What will be its value 
Reply fast

Grade:11

4 Answers

G.VENKATESH
7 years ago
The answer for the given question value of cos(-660)*tan(510) is  ““-0.2886”“.Thanks for asking this Question….….…
Poosala Vedaprakash
7 years ago
the answer of the question is -0.288675135
thanks for asking this question i will do my level best i feel this is good question
Ajay
7 years ago
COS (-660) * Tan (510)
= Cos (660) * Tan (510) Since cos (-x) = Cos x
= Cos (360 + 300 ) * Tan (360 + 150 )
 = Cos 300 * Tan 150
= Cos (360 – 60) * Tan (180-30)
= Cos 60 * (-Tan 30 )
= ½ * (-1/sqrt 3)
= –1/2 * 1/ sqrt 3
 
raj
7 years ago
hello,
Cos (660) * Tan (510) Since cos (-x) = Cos x
= Cos (360 + 300 ) * Tan (360 + 150 )
 = Cos 300 * Tan 150
= Cos (360 – 60) * Tan (180-30)
= Cos 60 * (-Tan 30 )
= ½ * (-1/sqrt 3)
= –1/2 * 1/ sqrt 3